Demand schedule P=610-3Q
Supply schedule P= 20+2Q
Calculate price elasticity
1
Expert's answer
2021-08-19T12:24:43-0400
"Q_D=\frac{610}{3}-\frac{1}{3}p"

"610-3Q=20+2Q"

"5Q=590"

"Q=118"

"p=256"

"E=-\frac{1}{3}\times\frac{256}{118}=-0.72"